OpenStudy (anonymous):

I WILL FAN OR MEDAL! What value of b makes the statement true? 17 = b – (–6.2) A. –23.2 B. –10.8 C. 10.8 D. 23.2

OpenStudy (here_to_help15):

its C

OpenStudy (anonymous):

Thank you could you help with a couple of more?

OpenStudy (here_to_help15):

um sure why not

OpenStudy (here_to_help15):

?

OpenStudy (anonymous):

Sam had a full 1 L bottle of olive oil. He poured 0.25 L of the olive oil into a container to make salad dressing. He also poured an unknown amount x onto a plate as a bread dip. There was 0.6 L of olive oil remaining in the bottle. Which equation can be used to solve for x? A. 1 + 0.25 – x = 0.6 B. 1 + 0.25 + x = 0.6 C. 1 – 0.25 – x = 0.6 D. 1 – 0.25 + x = 0.6

OpenStudy (here_to_help15):

C. 1 – 0.25 – x = 0.6

OpenStudy (anonymous):

Thank you!

OpenStudy (here_to_help15):

np